# Algorithms Trained on Normal Chest X-rays Can Predict Health Insurance Types

### v1

@@ Abstract
Artificial intelligence is revealing what medicine never intended to encode. Deep vision models, trained on chest X-rays, can now detect not only disease but also invisible traces of social inequality. In this study, we show that state-of-the-art architectures (DenseNet121, SwinV2-B*, MedMamba) can predict a patient’s health insurance type, a strong proxy for socioeconomic status, from normal chest X-rays with significant accuracy (AUC $\approx$ 0.67 on MIMIC-CXR-JPG, 0.68 on CheXpert). The signal persists even when age, race, and sex are controlled for, and remains detectable when the model is trained exclusively on a single racial group. Patch-based occlusion reveals that the signal is diffuse rather than localized, embedded in the upper and mid-thoracic regions. This suggests that deep networks may be internalizing subtle traces of clinical environments, equipment differences, or care pathways; learning socioeconomic segregation itself. These findings challenge the assumption that medical images are neutral biological data. By uncovering how models perceive and exploit these hidden social signatures, this work reframes fairness in medical AI: the goal is no longer only to balance datasets or adjust thresholds, but to interrogate and disentangle the social fingerprints embedded in clinical data itself.
@@ 3.1 Health Insurance Prediction from CXRs
Results As Table 1 indicates, in MIMIC-CXR-JPG, all trained health insurance predicting models could achieve an AUC performance around 0.65, with MedMamba having the best average AUC (0.6669) on the insurance predicting task and SwinTransformer performed the worst (0.6261). On the other hand, the models trained by on CheXpert also had a better-than-random performance on the health insurance prediction, as the best performance is 0.6761. Notice that CheXpert contains less and smaller sized images compared to MIMIC-CXR-JPG, which may be the major cause of the AUC decrease in SwinTransformer and MedMamba. The result suggested that the models are not simply guessing and that chest X-rays contain certain insurance information for the model to learn from.
@@
tab:mainAUC DenseNet Swin MedMamba MIMIC 0.6385 0.6261 0.6669 CheXpert 0.6761 0.6168 0.6188
@@ 3.2 Localization of insurance information on Xray - Patch-based training
Results Figure 1 reveals the AUC performance on the insurance type prediction task for the Remove-One-Patch and Keep-One-Patch experiement. The lowest performance (0.6332) occurs when we remove the center part of the image, while the highest performance (0.6572) appears when pixels at the mid-lower corner are set to zero. Conversely, the highest performance (0.6541) in Keep-one-patch experiment lands in the top-left corner when it is the only part that is preserved in an image. AUC performance drops to the lowest (0.6063) when only the bottom-right part of the image is retained. While the pulmonary area is located at the middle of a chest radiograph, these results suggest that the health insurance information is not only inferred from lungs. Combining the two experiments, the upper two-third part seems to contain the most health insurance type information, while the bottom part of a chest radiograph contains the least of it.
@@ 3.3 Experiments on Demographic Mediators
Result Table 2 shows the performance of multiple machine learning models to predict health insurance from the combination of age, race, and sex attributes. Among the five machine learning models (Random Forest, XGBoost, CatBoost, Decision Tree, KNN) that we have evaluated, XGBoost have the highest AUC value, but could only reach 0.5905. Therefore, we concluded that the results do not indicate a strong correlation between health insurance types and the three demographic features, as all the AUCs are no higher than 0.6. Table 3 demonstrates the health insurance type prediction performance of the DenseNet121 model trained solely from White people in our dataset. We observed that the performance is relatively consistent with training our model with the entire dataset, except a noticeable drop DenseNet. In this experiment, we could still observe the superiority of MedMamba over the other two models. Our results indicate that the ability of deep vision models to predict health insurance types is not mediated by race features.
@@
tab:race_stratified_AUC DenseNet Swin MedMamba 0.6047 0.6290 0.6679
@@ Appendix B Model architecture
All of the experiments utilized Lion lion as their optimizer, and the learning rate was 1e-6 for DenseNet, 1e-5 for SwinTransformer and MedMamba. The learning rate of DenseNet was adjusted to 1e-5 in experiment 3.3 for better performance. The batch size was 32 across our experiments, and we did not implement dropout or weight decay for regularization for simplicity. In MIMIC-CXR-JPG, all images were resized to 448x448 otherwise mentioned, while in CheXpert, all images were resized to 320x320 since the original image size in CheXpert is smaller than in MIMIC-CXR-JPG. All images were preprocessed with RandomHorizontalFlip, RandomRotation and Normalization in PyTorch for augmentation. A cross-entropy loss was utilized for health insurance type classification tasks. We trained our model on the training dataset, and selected the weights by the best performing average area under receiver operating characteristic curve (AUC) on the validation dataset. The final performance was tested on the testing dataset with the chosen weights. All experiments were trained on a single NVIDIA GeForce RTX 3090 or RTX 6000 Ada GPU.

### v3

@@ Abstract
Artificial intelligence is revealing what medicine never intended to encode. Deep vision models, trained on chest X-rays, can now detect not only disease but also invisible traces of social inequality. In this study, we show that state-of-the-art architectures (DenseNet121, SwinV2-B*, MedMamba) can predict a patient’s health insurance type, a strong proxy for socioeconomic status, from normal chest X-rays with significant accuracy (AUC $\approx$ 0.70 on MIMIC-CXR-JPG, 0.68 on CheXpert). The signal persists even when age, race, and sex are controlled for, and remains detectable when the model is trained exclusively on a single racial group. Patch-based occlusion reveals that the signal is diffuse rather than localized, embedded in the upper and mid-thoracic regions. This suggests that deep networks may be internalizing subtle traces of clinical environments, equipment differences, or care pathways; learning socioeconomic segregation itself. These findings challenge the assumption that medical images are neutral biological data. By uncovering how models perceive and exploit these hidden social signatures, this work reframes fairness in medical AI: the goal is no longer only to balance datasets or adjust thresholds, but to interrogate and disentangle the social fingerprints embedded in clinical data itself.
@@ 3.1 Health Insurance Prediction from CXRs
Results As Table 1 indicates, in MIMIC-CXR-JPG, all trained health insurance predicting models could achieve an AUC performance around 0.65, with DenseNet having the best average AUC (0.7007) on the insurance predicting task and SwinTransformer performed the worst (0.6182). On the other hand, the models trained by on CheXpert also had a better-than-random performance on the health insurance prediction, as the best performance is 0.6802. Notice that CheXpert contains less and smaller sized images compared to MIMIC-CXR-JPG, which may be the major cause of the AUC decrease in all models. The result suggested that the models are not simply guessing and that chest X-rays contain certain insurance information for the model to learn from.
@@
tab:mainAUC DenseNet Swin MedMamba MIMIC 0.7007 0.6182 0.6684 CheXpert 0.6802 0.6063 0.6183
@@ 3.2 Localization of insurance information on Xray - Patch-based training
Results Figure 1 reveals the AUC performance on the insurance type prediction task for the Remove-One-Patch and Keep-One-Patch experiement. The lowest performance (0.6386) occurs when we remove the center part of the image, while the highest performance (0.6624) appears when pixels at the bottom-right corner are set to zero. Conversely, the highest performance (0.6576) in Keep-one-patch experiment lands in the top-left corner when it is the only part that is preserved in an image. AUC performance drops to the lowest (0.6093) when only the mid-bottom part of the image is retained. While the pulmonary area is located at the middle of a chest radiograph, these results suggest that the health insurance information is not only inferred from lungs. Combining the two experiments, the upper two-third part seems to contain the most health insurance type information, while the bottom part of a chest radiograph contains the least of it.
@@ 3.3 Experiments on Demographic Mediators
Result Table 2 shows the performance of multiple machine learning models to predict health insurance from the combination of age, race, and sex attributes. Among the five machine learning models (Random Forest, XGBoost, CatBoost, Decision Tree, KNN) that we have evaluated, XGBoost have the highest AUC value, but could only reach 0.5905. Therefore, we concluded that the results do not indicate a strong correlation between health insurance types and the three demographic features, as all the AUCs are no higher than 0.6. Table 3 demonstrates the health insurance type prediction performance of the MedMamba model trained solely from White people in our dataset. We observed that the performance is relatively consistent with training our model with the entire dataset. In this experiment, we could still observe the superiority of DenseNet over the other two models. Our results indicate that the ability of deep vision models to predict health insurance types is not mediated by race features.
@@
tab:race_stratified_AUC DenseNet Swin MedMamba 0.6954 0.6154 0.6714
@@ Appendix B Model architecture
All of the experiments utilized Lion lion as their optimizer, and the learning rate was 5e-6 for DenseNet, 1e-5 for SwinTransformer and MedMamba. The learning rate of DenseNet was adjusted to 1e-5 in experiment 3.3 for better performance. The batch size was 32 across our experiments, and we did not implement dropout or weight decay for regularization for simplicity. In MIMIC-CXR-JPG, all images were resized to 448x448 otherwise mentioned, while in CheXpert, all images were resized to 320x320 since the original image size in CheXpert is smaller than in MIMIC-CXR-JPG. All images were preprocessed with RandomHorizontalFlip, RandomRotation and Normalization in PyTorch for augmentation. A cross-entropy loss was utilized for health insurance type classification tasks. We trained our model on the training dataset, and selected the weights by the best performing average area under receiver operating characteristic curve (AUC) on the validation dataset. The final performance was tested on the testing dataset with the chosen weights. All experiments were trained on a single NVIDIA GeForce RTX 3090 or RTX 6000 Ada GPU.